OpenStudy (anonymous):

Power Series. Prove that sum (n/(n+1)!) = 1

OpenStudy (anonymous):

\[\sum_{0}^{\infty} (n/(n+1)!) = 1\]

OpenStudy (zarkon):

I'll give a hint... \[\frac{n}{(n+1)!}=\frac{n+1-1}{(n+1)!}=\frac{1}{n!}-\frac{1}{(n+1)!}\]

OpenStudy (anonymous):

há ! Good one ! THANK YOU ! hahahahaha brilhant

OpenStudy (zarkon):

:)